# Subsampled Ensemble Can Improve Generalization Tail Exponentially

### v3

@@ Abstract
Ensemble learning is a popular technique to improve the accuracy of machine learning models. It hinges on the rationale that aggregating multiple weak models can lead to better models with lower variance and hence higher stability, especially for discontinuous base learners. In this paper, we provide a new perspective on ensembling. By selecting the best model trained on subsamples via majority voting, we can attain exponentially decaying tails for the excess risk, even if the base learner suffers from slow (i.e., polynomial) decay rates. This tail enhancement power of ensembling is agnostic to the underlying base learner and is stronger than variance reduction in the sense of exhibiting rate improvement. We demonstrate how our ensemble methods can substantially improve out-of-sample performances in a range of examples involving heavy-tailed data or intrinsically slow rates. Code for the proposed methods is available at https://github.com/mickeyhqian/VoteEnsemble.
@@ 1 Introduction
Ensemble learning (Dietterich, 2000; Zhou, 2012) is a class of methods to improve the accuracy of machine learning models. It comprises repeated training of models (the “base learners”), which are then aggregated through averaging or majority vote. In the literature, the main justification for ensemble methods, such as bootstrap aggregating (bagging) (Breiman, 1996) and boosting (Freund et al., 1996), pertains to bias/variance reduction or higher stability. This justification has been shown to be particularly relevant for certain U-statistics (Buja & Stuetzle, 2006) and models with hard-thresholding rules such as decision trees (Breiman, 2001; Drucker & Cortes, 1995).
Contrary to the established understanding, in this paper we present a new view of ensembling in offering an arguably stronger power than variance reduction: By suitably selecting the best base learners trained on random subsamples, ensembling leads to exponentially decaying excess risk tails. In particular, for general stochastic optimization problems that suffer from a slow, namely polynomial, decay in excess risk tails, ensembling can reduce these tails to an exponential decay rate. Thus, instead of the typical constant factor of improvement exhibited by variance reduction, our ensemble method offers a rate improvement, and moreover, the improvement is substantial.
In the following, we will first qualify our claims above by discussing how slow convergence can arise generically in machine learning and more general data-driven decision-making problems under heavy-tailed data. We then give intuition on our new ensembling perspective, proposed procedures, and the technicality involved in a full analysis.
Main results at a high level.
We begin by introducing a generic stochastic optimization problem
@@
where $\theta$ is the decision variable on space $\Theta$, $z\in\mathcal{Z}$ denotes the randomness governed by a probability distribution, and $l$ is the cost function. $n$ i.i.d. samples $\{z_{1},\ldots,z_{n}\}$ are available from the underlying distribution of $z$. In machine learning, $\theta$ corresponds to model parameters, $\{z_{1},\ldots,z_{n}\}$ the training data, $l$ the loss function, and $L$ the population-level expected loss. More generally, (1) encapsulates data-driven decision-making problems, namely the integration of data on $z$ into a downstream optimization task with overall cost function $l$ and prescriptive decision $\theta$. These problems are increasingly prevalent in various industrial applications (Kamble et al., 2020; Bertsimas et al., 2023; Ghosal et al., 2024), such as in supply chain network design where $\theta$ may represent the decision to open processing facilities, $z$ the uncertain supply and demand, and $l$ the total cost of processing and transportation.
@@
Given the data, we can train the model or decision with a learning algorithm that maps the data to an element in $\Theta$. This encompasses a wide range of methods, including machine learning training algorithms and data-driven approaches like sample average approximation (SAA) (Shapiro et al., 2021) and distributionally robust optimization (DRO) (Mohajerin Esfahani & Kuhn, 2018) in stochastic optimization. Our proposal and theory described below are agnostic to the choice of learning algorithm.
@@
We characterize the generalization performance of a solution to (1), denoted by $\hat{\theta}$, via the tail probability bound on the excess risk or regret $L(\hat{\theta})-\min_{\theta\in\Theta}L(\theta)$, i.e., $\mathbb{P}(L(\hat{\theta})>\min_{\theta\in\Theta}L(\theta)+\delta)$ for some fixed $\delta>0$, where the probability is over both the data and training randomness. By a polynomially decaying generalization tail, we mean that
$\mathbb{P}\Big{(}L(\hat{\theta})>\min_{\theta\in\Theta}L(\theta)+\delta\Big{)}\leq C_{1}n^{-\alpha}$ (2)
@@
for some $\alpha>0$ and $C_{1}$ depends on $\delta$. Such bounds are common under heavy-tailed data distributions (Kaňková & Houda, 2015; Jiang et al., 2020; Jiang & Li, 2021) due to slow concentration, which frequently arises in machine learning applications such as large language models (e.g, Jalalzai et al. (2020); Zhang et al. (2020); Cutkosky & Mehta (2021) among others), finance (Mainik et al., 2015; Gilli & Këllezi, 2006) and physics (Fortin & Clusel, 2015; Michel & Chave, 2007), and are proved to be tight (Catoni, 2012) for empirical risk minimization (ERM) (Vapnik, 1991). As our key insight, our proposed ensembling methodology can improve the above to an exponential decay, i.e.,
@@
where $k$ is the subsampled data size and can be chosen at a slower rate in $n$, and $\gamma<1$ depends on $k,\delta$ such that $\gamma\to 0$ as $k\to\infty$. Hence, when $k$ is properly chosen, the decay becomes exponential. This exponential acceleration is qualitatively different from the well-known variance reduction benefit of ensembling in several aspects. First, variance reduction refers to the smaller variability in predictions from models trained on independent data sets, which has a more direct impact on the expected regret than the tail decay rate. Second, the improvement by variance reduction is typical of a constant factor (e.g., Bühlmann & Yu (2002) reported a reduction factor of $3$), thus affecting at best the constant $C_{1}$ in (2), whereas we obtain an order-of-magnitude improvement.
Main intuition.
To facilitate our explanation, let us first focus on discrete space $\Theta$. Our ensembling methodology uses a majority-vote mechanism at the model level: After repeatedly running the learning algorithm on subsamples to generate many models, we output the model that occurs most frequently. This implicitly solves a surrogate optimization problem over the same decision space $\Theta$ as (1) that maximizes the probability of being output by the learning algorithm. This conversion of the original general objective in (1) to a probability objective is the key: As an expectation of a random indicator function, the latter is uniformly bounded even if the original objective is heavy-tailed. Together with a bootstrap argument that establishes the closeness between subsample and full data, this in turn results in exponentially decaying tails for the regret.
For more general problems with continuous space, we replace the simple majority vote with a vote based on the likelihood of being $\epsilon$-optimal among all the generated models when evaluated on a random subsample. This avoids the degeneracy issue of using a simple majority vote for continuous problems while retaining similar (in fact, even stronger as we will see) guarantees. Regardless of discrete or continuous model space, our main insight on turning (2) into (3) applies. Moreover, in the discrete case, it turns out that not only the tail bound but also the average-case regret improves exponentially. This also explains why our improvement is particularly significant for discrete-decision problems in the experiments.
The rest of the paper is organized as follows. Section 2 presents our ensemble methods and their finite-sample bounds. Section 3 presents experimental results, and Section 4 discusses related work. Section 5 discusses limitations and concludes the paper. A review of additional related work, technical proofs, and additional experimental results can be found in the appendix.
@@ 2 Methodology and theoretical guarantees
To solve (1) using data, we consider the generic learning algorithm in the form of a mapping
$\mathcal{A}(z_{1},\ldots,z_{n};\omega):\mathcal{Z}^{n}\times\Omega\to\Theta$
that takes in the training data $(z_{1},\ldots,z_{n})$ and outputs a model possibly under some algorithmic randomness $\omega$ that is independent of the data. Examples of $\omega$ include gradient sampling in stochastic first-order algorithms and feature/data subsampling in random forests. $\mathcal{A}(z_{1},\ldots,z_{n};\omega)$ serves as our base learner. For convenience, we omit $\omega$ to write $\mathcal{A}(z_{1},\ldots,z_{n})$ when no confusion arises.
@@ 2.1 A basic procedure
To understand $\mathsf{MoVE}$, we consider an optimization associated with the base learner $\mathcal{A}$
@@
which gives exactly the output of $\mathsf{MoVE}$. In other words, $\mathsf{MoVE}$ is a bootstrap approximation to the solution of (4). The following result materializes the intuition explained in the introduction on the conversion of the original potentially heavy-tailed problem (1) into a probability maximization (6) that possesses exponential bounds:
Theorem 1 (Finite-sample bound for Algorithm 1)
Consider discrete decision space $\Theta$. Recall $p_{k}(\theta)$ defined in (4). Let $p_{k}^{\max}:=\max_{\theta\in\Theta}p_{k}(\theta)$, $\mathcal{E}_{k,\delta}:=\mathbb{P}(L(\mathcal{A}(z_{1},\ldots,z_{k}))>\min_{\theta\in\Theta}L(\theta)+\delta)$ be the excess risk tail of $\mathcal{A}$, and
@@
For every $k\leq n$ and $\delta\geq 0$ such that $\eta_{k,\delta}>0$, the solution output by $\mathsf{MoVE}$ satisfies that
$+\exp\left(-\dfrac{B}{24}\cdot\dfrac{\eta_{k,\delta}^{2}}{\min\left(p_{k}^{\max},1-p_{k}^{\max}\right)+3\eta_{k,\delta}/4}\right)$
$+\mathbbm{1}\left(p_{k}^{\max}+\dfrac{\eta_{k,\delta}}{4}\leq 1\right)\cdot\exp\Big{(}-\dfrac{n}{2k}\cdot D_{\operatorname{KL}}\left(p_{k}^{\max}+\dfrac{\eta_{k,\delta}}{4}\Big{\|}p_{k}^{\max}\right)-\dfrac{B}{24}\cdot\dfrac{\eta_{k,\delta}^{2}}{1-p_{k}^{\max}+\eta_{k,\delta}/4}\Big{)}\Bigg{]}.$
In particular, if $\eta_{k,\delta}>4/5$, (8) is further bounded by
$\left|\Theta\right|\left(3\min\left(e^{-2/5},C_{1}\max(1-p_{k}^{\max},\;\mathcal{E}_{k,\delta})\right)^{\frac{n}{C_{2}k}}+e^{-B/C_{3}}\right),$ (9)
@@
Theorem 1 states that the excess risk tail of $\mathsf{MoVE}$ decays exponentially in the ratio $n/k$ and ensemble size $B$. The bound consists of three parts. The first part has two terms with the Kullback–Leibler (KL) divergences and arises from the bootstrap approximation of (4) with (5). The second part quantifies the Monte Carlo error in approximating (5) with a finite $B$. The third part comes from the interaction between the two sources of errors and is typically of higher order. The multiplier $\left|\Theta\right|$ in the bound is avoidable, e.g., via a space reduction as in our next algorithm.
@@
The quantity $\eta_{k,\delta}$ plays two roles. First, it quantifies how suboptimality in the surrogate problem (4) propagates to the original problem (1) in that every $\eta_{k,\delta}$-optimal solution for (4) is $\delta$-optimal for (1). Second, $\eta_{k,\delta}$ is directly related to the excess risk tail $\mathcal{E}_{k,\delta}$ of the base learner, in addition to $p_{k}^{\max}$ that captures the concentration of the base learner on $\delta$-optimal solutions. Therefore, $\eta_{k,\delta}$ taking large values signals the situation where the base learner already generalizes well. In this case, (8) can be simplified to (9). The bound (9) suggests that our approach does not hurt the performance of an already high-performing base learner as its generalization power is inherited through the $\max(1-p_{k}^{\max},\;\mathcal{E}_{k,\delta})$ term in the bound. See Appendix B for a more detailed comparison.
@@
The quantity $\eta_{k,\delta}$ also hints at how to choose the subsample size $k$. As long as $\eta_{k,\delta}$ is bounded away from 00, our bound decays exponentially fast. Therefore, $k$ can be chosen in such a way that the base learner outputs good models more often than bad ones in order for the exponential decay of our bound to take effect, but at the same time considerably smaller than $n$ to ensure the amount of acceleration. In the experiments, we choose $k=\max(10,n/200)$.
On the choice of $B$, note that the two KL divergences in the first part of the tail bound (8) are in general bounded below by $\mathcal{O}(\eta_{k,\delta}^{2})$ and so is the $\eta_{k,\delta}^{2}/(\min\left(p_{k}^{\max},1-p_{k}^{\max}\right)+3\eta_{k,\delta}/4)$ in the second part as $\eta_{k,\delta}$ is no larger than $1$. Therefore using an ensemble size of $B=\mathcal{O}(n/k)$ is sufficient to control the Monte Carlo error to a similar magnitude as the data error.

### v4

@@ Abstract
Ensemble learning is a popular technique to improve the accuracy of machine learning models. It traditionally hinges on the rationale that aggregating multiple weak models can lead to better models with lower variance and hence higher stability, especially for discontinuous base learners. In this paper, we provide a new perspective on ensembling. By selecting the best model trained on subsamples via majority voting, we can attain exponentially decaying tails for the excess risk, even if the base learner suffers from slow (i.e., polynomial) decay rates. This tail enhancement power of ensembling is agnostic to the underlying base learner and is stronger than variance reduction in the sense of exhibiting rate improvement. We demonstrate how our ensemble methods can substantially improve out-of-sample performances in a range of numerical examples involving heavy-tailed data or intrinsically slow rates. Code for the proposed methods is available at https://github.com/mickeyhqian/VoteEnsemble.
@@ 1 Introduction
Ensemble learning (Dietterich, 2000; Zhou, 2012) is a class of methods designed to improve the accuracy of machine learning models by combining multiple models, known as “base learners”, through aggregation techniques such as averaging or majority voting. In the existing literature, ensemble methods—most notably bagging (Breiman, 1996) and boosting (Freund et al., 1996)—are primarily justified based on their ability to reduce bias and variance or improve model stability. This justification has been shown to be particularly relevant for certain U-statistics (Buja & Stuetzle, 2006) and models with hard-thresholding rules, such as decision trees (Breiman, 2001; Drucker & Cortes, 1995).
In contrast to this traditional understanding, we present a novel perspective on ensemble learning, demonstrating its capability to achieve a significantly stronger effect than variance reduction: By suitably selecting the best base learners trained on random subsamples, ensembling leads to exponentially decaying excess risk tails. Specifically, for general stochastic optimization problems that suffer from a slow (polynomial) decay in excess risk tails, ensembling can reduce these tails to an exponential decay rate, a substantial improvement beyond the constant-factor gains typically exhibited by variance reduction.
To detail our contribution, we consider the generic stochastic optimization problem
@@
where $\theta\in\Theta$ is the decision variable, $z\in\mathcal{Z}$ is a random variable governed by some probability distribution, and $l(\cdot,\cdot)$ is the cost function. A dataset of $n$ i.i.d. samples $\{z_{1},\ldots,z_{n}\}$ is drawn from the underlying distribution of $z$. In the context of machine learning, $\theta$ corresponds to the model parameters, $\{z_{1},\ldots,z_{n}\}$ represents the training data, $l$ denotes the loss function, and $L$ is the population-level expected loss. More generally, (1) also encompasses data-driven decision-making problems, i.e., the integration of data on $z$ into a downstream optimization task with overall cost function $l$ and prescriptive decision $\theta$. These problems are increasingly prevalent in various industrial applications (Kamble et al., 2020; Bertsimas et al., 2023; Ghosal et al., 2024). For example, in supply chain network design, $\theta$ may represent the decision to open processing facilities, $z$ the uncertain supply and demand, and $l$ the total cost of processing and transportation.
@@
Given the data, a learning algorithm can be used to map the data to an element in $\Theta$, yielding a trained model or decision. This encompasses a variety of methods, including machine learning training algorithms and data-driven approaches such as sample average approximation (SAA) (Shapiro et al., 2021) and distributionally robust optimization (DRO) (Mohajerin Esfahani & Kuhn, 2018) in stochastic optimization. The theoretical framework and methodology proposed in this paper are agnostic to the choice of learning algorithms.
@@
Main Results at a High Level.    Let $\hat{\theta}$ be the output of a learning algorithm. We characterize its generalization performance through the tail probability bound on the excess risk $L(\hat{\theta})-\min_{\theta\in\Theta}L(\theta)$, i.e., $\mathbb{P}(L(\hat{\theta})>\min_{\theta\in\Theta}L(\theta)+\delta)$ for some fixed $\delta>0$, where the probability is over both the data and training randomness. A polynomially decaying generalization tail refers to:
$\mathbb{P}\Big{(}L(\hat{\theta})>\min_{\theta\in\Theta}L(\theta)+\delta\Big{)}\geq C_{1}n^{-\alpha},$ (2)
@@
for some $\alpha>0$ and $C_{1}$ that depends on $\delta$. Such bounds are common under heavy-tailed data distributions (Kaňková & Houda, 2015; Jiang et al., 2020; Jiang & Li, 2021) due to slow concentration, which frequently arises in machine learning applications such as large language models (Jalalzai et al., 2020; Zhang et al., 2020; Cutkosky & Mehta, 2021), finance (Mainik et al., 2015; Gilli & Këllezi, 2006), and physics (Fortin & Clusel, 2015; Michel & Chave, 2007). This can be illustrated with a simple linear program (LP):
Example 1.1 (LP with a polynomial tail).
Consider the stochastic LP $\min_{\theta\in[0,1]}\mathbb{E}[z\theta]$, i.e., $l(\theta,z)=z\theta$ and $\Theta=[0,1]$ in (1), and its SAA solution $\hat{\theta}\in\mathrm{argmin}_{\theta\in[0,1]}\sum_{i=1}^{n}z_{i}/n\cdot\theta$. Assume $z$ has a non-zero density everywhere and is symmetric with respect to its mean $\mathbb{E}\left[z\right]=1$ (hence $L(\theta)=\theta$). Then we have $\mathbb{P}(\hat{\theta}=1)\geq\mathbb{P}(\sum_{i=1}^{n}z_{i}/n<0)\geq\mathbb{P}(\sum_{i=1}^{n-1}z_{i}\leq n-1\text{ and }z_{n}<1-n)=\mathbb{P}(\sum_{i=1}^{n-1}z_{i}\leq n-1)\mathbb{P}(z_{n}<1-n)$, where the last equality uses the independence of $z_{i}$’s. By the symmetry of $z$, we have $\mathbb{P}(\sum_{i=1}^{n-1}z_{i}\leq n-1)=1/2$, thus the tail
$\mathbb{P}\Big{(}L(\hat{\theta})>\min_{\theta\in[0,1]}L(\theta)+\delta\Big{)}\geq\mathbb{P}(\hat{\theta}=1)\geq\frac{1}{2}\mathbb{P}(z<1-n).$ (3)
If $z$ has a polynomial tail, e.g., $\mathbb{P}(z<1-n)=\Omega(n^{-\alpha})$ for some $\alpha>0$ where $\Omega(\cdot)$ contains some multiplicative constant, the generalization tail bound (2) applies. Appendix E provides another example from linear regression.
As the key contribution of our work, we propose ensemble methods that significantly improve these bounds, achieving an exponentially decaying generalization tail:
@@
where $k$ is the subsample size and $\gamma<1$ depends on $k,\delta$ with $\gamma\to 0$ as $k\to\infty$. By appropriately choosing $k$ at a slower rate in $n$, the decay becomes exponential. This exponential acceleration is fundamentally different from the well-known variance reduction benefit of ensembling in two perspectives. First, variance reduction refers to the smaller variability in predictions from models trained on independent data sets, thus has a more direct impact on the expected regret than the tail decay rate. Second, variance reduction typically yields a constant-factor improvement (e.g., Bühlmann & Yu (2002) report a reduction factor of $3$), whereas we obtain an order-of-magnitude improvement.
Main Intuition.    Consider first the discrete space $\Theta$. Our ensemble method employs a majority-vote mechanism at the model level: The learning algorithm is repeatedly run on subsamples to generate multiple models, and the model appearing most frequently is selected as the output. This process effectively estimates the mode of the sampling distribution of the learned model by subsampling, and thus is less susceptible to extreme data and training randomness that incurs the slow tail decay in (2). This mode estimation can be formalized via a surrogate optimization problem over the same decision space $\Theta$ as (1) that maximizes the probability of being output by the learning algorithm. The probability objective, as the expected value of a random indicator function, is uniformly bounded even if the original objective is heavy-tailed and hence admits exponential decay in the tail. To illustrate the main idea using Example 1.1, although $\mathbb{P}(\hat{\theta}=1)$ can be substantial in a heavy-tailed setting, it holds that $\mathbb{P}(\hat{\theta}=0)>\mathbb{P}(\hat{\theta}=1)$, and thus the mode of $\hat{\theta}$, i.e., 00, recovers the optimal solution.
For general problems with continuous decision spaces, we replace the majority vote with a voting mechanism based on the likelihood of being $\epsilon$-optimal among all models when evaluated on random subsamples. This avoids the degeneracy of using a majority vote for continuous problems while retaining similar (in fact, even stronger) theoretical guarantees. For both discrete and continuous problems, our method fundamentally improves the tail behavior from (2) to (4).
Organization.    The rest of the paper is organized as follows. Section 2 presents our methods and their finite-sample bounds. Section 3 presents experimental results, Section 4 discusses related work, and Section 5 concludes the paper. A review of additional related work, technical proofs, and additional experiments can be found in the appendix.
@@ 2 Methodology and Theoretical Guarantees
We consider the generic learning algorithm in the form of
$\mathcal{A}(z_{1},\ldots,z_{n};\omega):\mathcal{Z}^{n}\times\mathbf{\Omega}\to\Theta$
that takes in the training data $(z_{1},\ldots,z_{n})\in\mathcal{Z}^{n}$ and outputs a model possibly under some algorithmic randomness $\omega\in\mathbf{\Omega}$ that is independent of the data. Examples of $\omega$ include gradient sampling in stochastic algorithms and feature/data subsampling in random forests. For convenience, we omit $\omega$ to write $\mathcal{A}(z_{1},\ldots,z_{n})$ when no confusion arises.
@@ 2.1 A Basic Procedure
To understand $\mathsf{MoVE}$ from the lens of mode estimation, we consider an optimization associated with the base learner $\mathcal{A}$
@@
which gives exactly the output of $\mathsf{MoVE}$. In other words, $\mathsf{MoVE}$ is a bootstrap approximation to the solution of (5). The following result materializes the intuition explained in Section 1 on the conversion of the original potentially heavy-tailed problem (1) into a probability maximization (7) that possesses exponential bounds.
Theorem 2.1 (Informal bound for Algorithm 1).
Consider discrete decision space $\Theta$. Let $p_{k}^{\max}:=\max_{\theta\in\Theta}p_{k}(\theta)$, where $p_{k}(\theta)$ is defined in (5). For any $\delta>0$, denote $\mathcal{E}_{k,\delta}:=\mathbb{P}(L(\mathcal{A}(z_{1},\ldots,z_{k}))>\min_{\theta\in\Theta}L(\theta)+\delta)$ as the excess risk tail of $\mathcal{A}$, and
@@
Then, for every $k\leq n$ and $\delta\geq 0$ such that $\eta_{k,\delta}>0$, the solution output by $\mathsf{MoVE}$ satisfies that
$\ \mathbb{P}(L(\hat{\theta}_{n})>\min_{\theta\in\Theta}L(\theta)+\delta)$ (9)
$\leq$ $\left|\Theta\right|\Big{[}4\exp(-\Omega(n/k))+\exp(-\Omega(B))\Big{]},$
where $\left|\Theta\right|$ denotes the cardinality of $\Theta$, and $\Omega(\cdot)$ contains multiplicative coefficients that depend on $p_{k}^{\max}$ and $\eta_{k,\delta}$. If $\eta_{k,\delta}>4/5$, (9) is further bounded by
$\left|\Theta\right|\Big{(}3\min\big{\{}e^{-2/5},C_{1}\max\left\{1-p_{k}^{\max},\mathcal{E}_{k,\delta}\right\}\big{\}}^{\frac{n}{C_{2}k}}+e^{-\frac{B}{C_{3}}}\Big{)},$ (10)
where $C_{1},C_{2},C_{3}>0$ are universal constants.
@@
The formal statement of Theorem 2.1 is deferred to Theorem C.7 in Appendix C.2. Theorem 2.1 states that the excess risk tail of $\mathsf{MoVE}$ decays exponentially in the ratio $n/k$ and ensemble size $B$. The bound (9) consists of two terms: The term $\exp(-\Omega(n/k))$ arises from the bootstrap approximation of (5) with (6), whereas the term $\exp(-\Omega(B))$ quantifies the Monte Carlo error in approximating (6) with a finite $B$. The multiplier $\left|\Theta\right|$ in the bound is avoidable, e.g., via a space reduction as in our next algorithm.
@@
The quantity $\eta_{k,\delta}$ plays two roles. First, it quantifies how suboptimality in the surrogate problem (5) propagates to the original problem (1) in that every $\eta_{k,\delta}$-optimal solution for (5) is $\delta$-optimal for (1). Second, $\eta_{k,\delta}$ is directly related to the excess risk tail $\mathcal{E}_{k,\delta}$ of the base learner, in addition to $p_{k}^{\max}$ that captures the concentration of the base learner on near-optimal solutions. Therefore, a large $\eta_{k,\delta}$ signals the situation where the base learner already generalizes well. In this case, (9) reduces to (10). (10) suggests that our approach does not hurt the performance of an already high-performing base learner as its generalization power is inherited through the $\max\left\{1-p_{k}^{\max},\;\mathcal{E}_{k,\delta}\right\}$ term in the bound. See Appendix B for a more detailed discussion.
@@
Theorem 2.1 also hints at how to choose the hyperparameters, $k$ and $B$. As long as $\eta_{k,\delta}$ is bounded away from 00, our bound decays exponentially fast. Therefore, $k$ can be chosen such that the base learner outputs good models more often than bad ones in order for the exponential decay to take effect, but at the same time considerably smaller than $n$ to make the bound (9) meaningful. In the experiments, we choose $k=\max(10,n/200)$. Regarding the choice of the ensemble size $B$, we observe from (9) that using a $B=\mathcal{O}(n/k)$ is sufficient to control the Monte Carlo error to a similar magnitude as the statistical error.
Applying Theorem 2.1 to Example 1.1 gives an exponential tail as opposed to the original slow decay in (3).
Corollary 2.2 (Enhanced tail for Example 1.1).
Consider the stochastic LP in Example 1.1 and denote $q_{k}:=\mathbb{P}(\sum_{i=1}^{k}z_{i}>0)$. We have $q_{k}>1/2$ by the symmetry of $z$. If $\mathsf{MoVE}$ is applied to Example 1.1 with $\mathcal{A}$ being the SAA, we have $p_{k}^{\max}=q_{k}$, $\eta_{k,\delta}=2q_{k}-1$ whenever $\delta<1$, and $\left|\Theta\right|=2$. Consequently, $\eta_{k,\delta}>0$ for every $k>0$ and $\delta<1$, ensuring the tail upper bound (9) holds. If additionally $q_{k}>0.9$, we also have the tail upper bound $6\min\big{\{}e^{-2/5},C_{1}(1-q_{k})\big{\}}^{\frac{n}{C_{2}k}}+2e^{-\frac{B}{C_{3}}}$ from (10).
The proof of Corollary 2.2 can be found in Appendix C.3.